The series combination of five cells, each electromotive force of 1.5 V and internal resistance of 0.5Ω, is connected to an external resistance of 5.0 Ω. Find the (a) total internal resistance, (b) electromotive force of the battery, (c) total resistance of the circuit, (d) current through the external circuit, (e) terminal voltage of the battery, and (f) terminal voltage of a cell.

1
Expert's answer
2022-02-16T10:27:59-0500

a) Total internal resistance "=5\times 0.5=2.5 \Omega"

b) electromotive force of the battery = "1.5\times 5=7.5v"

c) total resistance of the circuit = 5+2.5=7.5"\Omega"

d) current through the external circuit "=V\/R=7.5\/7.5=1.0A"

e) terminal voltage of the battery, "E=IR-Ir=1\times 7.5-1\times 2.5=5V"

f) terminal voltage of a cell= 1.5V - "0.5\times1=1.0V"